OpenStudy (anonymous):

Derivative of f(x) = -3 x^5 sqrt[x] + 4/(x^2 sqrt[x])

OpenStudy (anonymous):

\[f(x) = -3x^5 \sqrt{x} + \frac{ 4 }{ x^2 \sqrt{x} }\]

OpenStudy (anonymous):

It looks tedious since it involves product and quotient rule. Anyway, what is the product and quotient rule formulas?

geerky42 (geerky42):

Actually, sum rule and power rule is only needed here. See: \(\large x^n\sqrt{x} = x^nx^{1/2} = x^{n+1/2}\)

geerky42 (geerky42):

For second term, just remember that \(\dfrac{1}{x^n} = x^{-n}\)

geerky42 (geerky42):

Sum Rule: \(\dfrac{d}{dx}[f(x)\pm g(x)] = f'(x)\pm g'(x)\) Power Rule: \(\dfrac{d}{dx}x^n = nx^{n-1}\)

geerky42 (geerky42):

@4586Jen Does that help?